what is the sum of:



5 2/6
+6 1/6
_______

Found 2 solutions by longjonsilver, lenny460:
Answer by longjonsilver(2297)

11 and 3/6 which is 11 and a half
jon

Answer by lenny460(1073)

5 2/6 + 6 1/6
5 + 6 = 11
2/6 + 1/6 = 3/6 = 1/2
11 + 1/2 = 11 1/2
Answer:
11 1/2